Please help.
When an employee produces x units per hour, the hourly wage in dollars is given
by y = 0.75x + 8. What are the hourly wages for the followig number of units
2, 3, 10, 15 and 20.
Thanks
JMS

Answer by checkley71(8403)

Y=.75X+8
Y=.75*2+8 Y=1.5+8 Y=9.5 ANSWER.
Y=.75*3+8 Y=2.25+8 Y=10.25
Y=.75*10+8 Y=7.5+8 Y=15.5
Y=.75*20+8 Y=15+8 Y=23